OpenStudy (anonymous):

Write the equation in its equivalent logarithmic form 2^3=x

OpenStudy (zzr0ck3r):

2^3 = x log_2(x) = 3

OpenStudy (unklerhaukus):

\[b^x=a\]\[\log_ba=x\]

OpenStudy (anonymous):

log2^x=3

OpenStudy (anonymous):

oh wow im late

OpenStudy (anonymous):

 like this

OpenStudy (anonymous):

the equation tool is whack... lol\[\log_2 x=3\]

OpenStudy (anonymous):

follow rhaukus' equation model.

OpenStudy (anonymous):

Thank you everyone